OpenStudy (anonymous):

what is the simplified form of square root 98n^19 A. 2n^9 square root 7 B. 7n^4 square root 2n^3 C. 7n^9 square root 2n D. 4n^9 square root 6n

OpenStudy (ash2326):

We have \[\large \sqrt{98n^{19}}\] Factors of 98 \[98= 1\times 2 \times 7 \times 7\] also \[n^{19}= n^9 \times n^9 \times n\] so \[\large \sqrt{ 2 \times 7 \times 7\times n^9 \times n^9 \times n}\] So we have \[ 7\times n^9 \sqrt{2 \times n}\] or \[\large 7n^9 \sqrt{2n}\]